OpenStudy (anonymous):

If the square of the length of the longest side of a triangle is greater than the sum of the squares of the lengths of the other two sides, then what type of triangle was formed?

OpenStudy (anonymous):

@wolf1728 its just one word ;3

OpenStudy (wolf1728):

I believe no triangle was formed. Basically, according to the triangle inequality theorem, http://1728.org/trianinq.htm in any triangle the longest side must be less than the sum of the other 2 sides. I'm making the assumption that the square of the longest side and the square of the sum of the other 2 sides will have the same relationship as the longest side compared to the sum of the other sides.

OpenStudy (wolf1728):

Okay in one word -NONE

OpenStudy (anonymous):

It's either A.Right B.obtuse C.Acute D.Scalene.............

OpenStudy (anonymous):

There is no answer maked "none"

OpenStudy (wolf1728):

Okay let's say we have a "triangle" of sides 25, 11 and 10 (which does NOT form a triangle). Squaring 25 = 625 11² + 10² = 121 + 100 Summing those squares = 221. There we have the square of the longest side being greater than the sum of the squares of the other 2 sides. Please tell me what kind of triangle is formed?

OpenStudy (anonymous):

it makes no sense AHAH

OpenStudy (wolf1728):

Ahhhhhhhhh so either • the question is phraed incorrectly or • they forgot to put an 'E' choice.

OpenStudy (anonymous):

Can you help with this one?

OpenStudy (wolf1728):

WX YX and YW I could explain why if you want.

OpenStudy (anonymous):

If the measure of two sides of a triangle are 3 feet and 7 feet, what is the least possible measure for the third side? Would it be 3 or 4?

OpenStudy (wolf1728):

Must greater than 4. (There's a calculator for that in the link I posted).

OpenStudy (anonymous):

Must be greater than 4?

jigglypuff314 (jigglypuff314):

7 + 3 = 10 7 - 3 = 4 therefore 4 < third side < 10

OpenStudy (anonymous):

so 5 :D

OpenStudy (wolf1728):

Yes - if it equals 4 you'd just have a straight line.

jigglypuff314 (jigglypuff314):

yeah that would work :)

jigglypuff314 (jigglypuff314):

11 + 5 = 16 11 - 5 = 6 so between 6 and 16

OpenStudy (wolf1728):

6 < third side < 16 Jiggly are you using that calculator?

OpenStudy (anonymous):

Would it b 6m<x<16m or 6m>x>16m

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Ik it's between 5.4 and 13 but which direction? idk

jigglypuff314 (jigglypuff314):

can't see it?

OpenStudy (anonymous):

5.4<x<13 Right?

jigglypuff314 (jigglypuff314):

if it's between 5.3 and 13 then yes :)

OpenStudy (anonymous):

Whoops there it is ;3

OpenStudy (anonymous):

And this one is weird. It should be "none" but there is no choice like that :/

OpenStudy (wolf1728):

5.4 < third side < 13

OpenStudy (anonymous):

Got it already! Thanks wolf!

jigglypuff314 (jigglypuff314):

that one is pythagorean theorem put into words :)

OpenStudy (anonymous):

But which one is itttt?

OpenStudy (anonymous):

No triangle could be formed :/

jigglypuff314 (jigglypuff314):

the pythagorean theorem can only be used for right triangles

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

#stupid

OpenStudy (anonymous):

Wolf can you give her a medal? ;3

OpenStudy (wolf1728):

And since the hypotenuse is greater than the sum of the squares of the other 2 sides then we have what kind of triangle?

OpenStudy (anonymous):

Obtuse

OpenStudy (anonymous):

@jigglypuff314 How much time do you have left? Before you leave?

OpenStudy (wolf1728):

I still think it is NO triangle.

jigglypuff314 (jigglypuff314):

well... I've still got my own homework to do so... asap kinda ;) If you still need me around I'll stay though :P and it would be a right triangle because it's talking about the sides "squared"

OpenStudy (anonymous):

The answer was obtuse lmfao......... i got the question wrong but I still got a 9/10 which is an A :D

OpenStudy (wolf1728):

I searched for an obtuse triangle and it is attached. Notice the longest side is smaller than the other sides.

OpenStudy (wolf1728):

That question is worded very badly. Let's suppose we have sides of 1, 1 and 100. Longest side squared is > 1² + 1² What kind of triangle does it form?